OpenStudy (anonymous):

Solve: 18/(s+v) = 18/(s-v) -9

OpenStudy (anonymous):

which one is unknown: s or v?

OpenStudy (anonymous):

both of them

OpenStudy (anonymous):

the ans seems to be 4v=s^2 - v^2 , but not sure how

OpenStudy (anonymous):

you need two equation to uniquely solve for two unknowns... possible answers: s = 0, v = 4, or s = 1, v = -2+/- sqrt(5). and so on and on...

OpenStudy (anonymous):

so now: which one is your unknown?

OpenStudy (anonymous):

Yes, this is the first one, confused with this one, the ans seems to be 4v=s^2 - v^2 , but not sure how

OpenStudy (anonymous):

the ans of the first equation

OpenStudy (anonymous):

let me show you guys how to do this: first you HAVE TO assume one variable is known. you CANNOT have two vars there in ONE eq and hope to solve it. let's say s is known, then you can easily convert this equation into a quadratic eq by multiplying both sides by (s-v)(s+v), then solve for v, and the answer is: v = +/-sqrt(s^2+4)-2. you can also do the same for s expressed with v.

OpenStudy (anonymous):

if you want integer solution.. then that is it: s = 0, and v= 4.

OpenStudy (anonymous):

otherwise, what you get is a Hyperbola

myininaya (myininaya):

\[\frac{18}{s+v}*(s-v)(s+v)=\frac{18}{s-v}*(s-v)(s+v)-9*(s+v)(s-v)\] \[18(s-v)=18(s+v)-9(s^2-v^2)\] \[18s-18v=18s+18v-9s^2+9v^2\] \[18s-18s-18v-18v=-9s^2+9v^2\] \[-36v=-9s^2+9v^2\] \[\frac{-36v}{9}=\frac{-9s^2}{9}+\frac{9v^2}{9}\] \[-4v=-s^2+v^2\] \[(-1)*-4v=(-1)(-s^2+v^2)\] \[4v=s^2+v^2\]

myininaya (myininaya):

but apple is right we need to know what to solve for

myininaya (myininaya):

it just say solve

OpenStudy (anonymous):

i did not say your answer is wrong, but i hope you also understand that rewriting an equation is not the same as solving an eq.... you have to move your unknowns to one side, and knowns to the other.. that is called solving an eq. even if it says solve for 4v, you still have v on the other side....

myininaya (myininaya):

yes i know lol

myininaya (myininaya):

i'm trying to say it would be nice for it to say solve for something instead of solve

OpenStudy (anonymous):

right. good luck.

myininaya (myininaya):

you too